OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals eight divided by x and g of x equals eight divided by x

OpenStudy (anonymous):

f(x)=8/x g(x)=8/x is that right?

OpenStudy (anonymous):

yes

OpenStudy (psymon):

Replace the x in f(x) with the whole function g(x) and then do the same with with g(x). If both times you can simplify the function down to equal x, then they are inverses.

OpenStudy (anonymous):

well then you just plug in f(g(x)) = x f(x)=8/x g(x)=8/x first lets do f f(8/x)=8/x = 8/(8/x)= 8* x/8 =x

OpenStudy (anonymous):

so that is is true and same goes for g g(8/x)=8/x= 8/(8/x) = x cuzz the 8 cancels

OpenStudy (anonymous):

thanks youu(: